OpenStudy (anonymous):

A) What is the slope of the line through these points? B) Write the equation in point slope form using the slope you found in part A and one of the ordered pairs given. It doesn’t matter which ordered pair you use. The point slope formula is (y – y1) = m(x – x1). C) Simplify your answer in part B and write the equation in slope-intercept form. Show all work.

OpenStudy (mef4.0):

what are the given points?

OpenStudy (anonymous):

(2,1) and (3,4), sorry

OpenStudy (mef4.0):

use the slope formula first: its the difference of the y's over the difference of the x's so, 4-1 over 3-2 or 3/1 so your slope is 3. That's your m in the point slope form

OpenStudy (mef4.0):

now take either point and sub in the values for y1 and x1 it does not matter which ordered pair you use - just don't get your x and y mixed up

OpenStudy (mef4.0):

to get it into slope-intercept form, you'll need to simplify that so that y is by itself and everything else is on the other side of the = sign. Does that make sense?

OpenStudy (anonymous):

would it be (1-1) = 3(3-2) this is saying that I'm using the (2,1) point

OpenStudy (mef4.0):

you substitute values for the x1 and y1 ONLY. You must leave in the x and y so you'll get: y-1 = 3(x-2) or y-4 = 3(x-3) either one will be correct for part b of your question

OpenStudy (anonymous):

and then how do I simplify that ?